# Ankle Exoskeletons May Hinder Standing Balance in Simple Models of Older and Younger Adults

### v2

@@ Abstract
Humans rely on ankle torque to maintain standing balance, particularly in the presence of small to moderate perturbations. Reductions in maximum torque (MT) production and maximum rate of torque development (MRTD) occur at the ankle with age, diminishing stability. Ankle exoskeletons are powered orthotic devices that may assist older adults by compensating for reduced muscle force and power production capabilities. They may also be able to assist with ankle strategies used for balance. However, no studies have investigated the effect of such devices on balance in older adults. Here, we model the effect ankle exoskeletons have on stability in physics-based models of healthy young and old adults, focusing on the mitigation of age-related deficits such as reduced MT and MRTD. We show that an ankle exoskeleton moderately reduces feasible stability boundaries in users who have full ankle strength. For individuals with age-related deficits, there is a trade-off. While exoskeletons augment stability in low velocity conditions, they reduce stability in some high velocity conditions. Our results suggest that well-established control strategies must still be experimentally validated in older adults.
@@ 1 Introduction
Ankle exoskeletons are powered orthotic devices that can assist people with mobility impairments, such as older adults, by compensating for reduced muscle force and power capabilities. Typically, they provide actuation about the ankle joint only in the sagittal plane, generating a push-off torque during dynamic activities such as walking and stair climbing. Exoskeletons have been shown to reduce metabolic cost of transport during walking, in both younger [1] and older adults [2]. The main source of their appeal to older adults, however, is their perceived potential to reduce fall risk [3]. It is therefore crucial to develop our understanding of the effect of ankle exoskeleton assistance on stability, which is currently incomplete.
@@
Torque production at the ankle is an important contributor to balance, which suggests that ankle exoskeletons have the potential to improve stability during standing. Unimpaired adults tend to rely on ankle torques to maintain a stable standing position, only switching to a hip strategy in response to large perturbations [4]. In older adults, maximum available plantar flexion and dorsiflexion torques are lower, reducing the range of perturbations that can be accommodated with an ankle strategy [5]. This reduction in peak ankle torque has been shown in multiple studies to be correlated with decreased performance in balancing tasks [5] and a higher risk of falls [6, 7]. The rate at which this torque can be produced also declines with age [8], although the effect of a lower maximum rate of torque development (MRTD) on standing balance is inconclusive. Reduced MRTD at the knee is associated with a history of falls [9, 10], while plantar flexor MRTD is significantly correlated with lower performance on a single standing leg balance test for older men, but not for older women [11]. Other studies have found that lower ankle MRTD is not a strong predictor of fall likelihood [12, 13]. Still, it is possible that lower MRTD contributes, independently of lower MT, to increased fall risk in older adults.
@@
Although ankle exoskeletons can, in theory, mitigate both MT and MRTD deficits by producing large amounts of torque quickly, the effect they have on stability is unclear. Depending on the population, device, controller, and perturbation paradigm, ankle exoskeletons have been reported to enhance stability [14, 15, 16], slightly reduce stability [17, 15], or have no effects on stability [18, 19]. Notably, the majority of balance studies for powered exoskeletons have focused on young and able bodied users, who do not have impaired stability. As these devices are likely to be marketed to enhance mobility in older populations [3], it is imperative to understand how these they may affect the stability of users with age-related reduction in MT and MRTD. We currently do not know how these age-related changes may interact with additional exoskeleton assistance at the ankle. Because perturbation experiments on older adults are expensive, time-consuming, and risky, developing a generalizable model-based characterization of standing balance with ankle exoskeletons is a useful approach. Modeling can provide insights into the mechanisms underlying exoskeleton-assisted standing balance, including how constraints on the human-exoskeleton system are affected both by the device and by age-related joint-level torque deficits.
@@
Bounds on human stability for feet-in-place activities such as standing may be characterized by computing the set of all body center of mass (CoM) positions and velocities from which it is possible to stabilize to quiet standing, meaning that the CoM position is above the foot and the CoM velocity is close to zero [20]. This set is sometimes called the feasible stability region [21, 22]. In the case of a simplified linear inverted pendulum these are the well-known extrapolated center of mass boundaries [23]. In this text we will use the term stabilizable region, abbreviated ‘SR,’ as in our previous work.
@@
Stabilizable regions have been computed for various legged dynamical systems [24, 25], including one example where the boundary was partially computed for a model of a human wearing an optimally controlled hip-knee-ankle exoskeleton [26]. The full stabilizable region under aging-related biological constraints and different ankle exoskeleton controllers has not been explored. Furthermore, formal stability properties of this region have not been defined. One particularly important attribute is invariance. If the SR is controlled invariant, this means that there always exists a controller that allows the human to stay within the region. If it is not, then there may be states within this region from which the human can pass through the SR, but not stabilize. Invariance guarantees that the SR does not include states which merely pass through the set. This property is particularly important to verify when there are constraints on the control input. Such constraints may correspond to biologically meaningful quantities such as maximum joint torque, or to environmental constraints such as the frictional forces on the support surface.
@@
In a previous paper, we introduced a general framework for determining the complete SR boundary for human-exoskeleton dynamical systems [27]. Our approach formulates the SR boundary as the solution of a single partial differential equation rather than an algorithmic search. In contrast to prior work, we include strong formal invariance and stability guarantees over the entire SR in the presence of both biological and environmental constraints.
@@
Here we provide a generalizable, model-based understanding of how an ankle exoskeleton may alter the domain of invariant stable motions available to a user in the presence of joint level functional changes associated with aging, including reduced MT production and MRTD. To our knowledge, this is the first study that attempts to systematically understand the combined effect of ankle exoskeleton assistance and age-related ankle torque deficits on feasible stability. Using common exoskeleton control strategies, we analyze the effect that exoskeletons have of the invariant stabilizable region for models of healthy young and healthy old adults. We show that while exoskeletons can indeed enhance ankle torque production and improve stability under certain conditions, they can also act as a disturbance, resulting in a trade-off between increased torque bounds and stability.
@@ 2 Preliminaries
Our methods for computing stabilizable regions (SRs) are based on two notions, controlled invariance and backward reachability, which we formally define in the following sections.
@@ 2.1 Controlled invariance and backward reachability
Figure 1: Illustration of a target set (green) and its backward reachable set (purple). The state $\hat{x}_{0}$ is in the BRS because there exists a controller $u_{1}$ that drives the system to the target set at time $T$. No such controller exists for $\tilde{x}_{0}$ so it is not in the BRS. Note that controller $u_{2}$ does not drive the system to the target. If the target set is controlled invariant then the union of the target set and the BRS is also invariant. Here, the BRS was computed with model parameters corresponding to a young male, using our method described in Section 4. The gray region is computed using the extrapolated center of mass (XCoM) method from [23], which we have converted here to angular coordinates.
@@
A set $\Omega\subset X$ is controlled invariant if for any initial condition in the set, there exists a controller that maintains the system state within $\Omega$:
@@
The BRS only contains states from which the system can reach the target set at precisely time $T$. The set of all states from which the target set can be reached at $t\leq T$ is called the backward reachable tube, and is a union of backward reachable sets:
@@ 2.2 Relating invariance and reachability to human stability
The relationship between human stability and our formal mathematical definitions is illustrated in Figure 1. Assume that the dynamics (1) represent a parameterized model of a standing human. In the figure, the green ellipse represents a controlled invariant target set of states corresponding to quiet standing. Note that for the purposes of the illustration the set is depicted to be larger than it is. The purple region is the true-to-size backward reachable set that we compute using a constrained dynamical model with parameters corresponding to a young male, as described later in Section 3. As labeled in the figure, the boundaries of this set correspond to forward and backward fall boundaries and to constraint failures. Outside of these boundaries the ankle strategy fails and the model must take some action to avoid a fall, such as taking a step. For visual clarity we only depict the portion of the state space corresponding to forward angular velocities. Because the target set is controlled invariant, the resulting purple safe region is as well. For comparison, we show the same bounds computed using the extrapolated center of mass method, from [23], in light gray.
To determine the stability bounds of the human-exoskeleton system, we also compute reachable sets with the dynamics of an ankle exoskeleton controller added to (1), the details of which are presented in Section 3 and 4. We construct invariant target sets corresponding to quiet standing with exoskeleton assistance, and then compute the backward reachable sets. The resulting sets then delineate the feasible stabilizable boundaries with ankle exoskeleton assistance.
@@ 3 Human-Exoskeleton Model
Figure 2: Free body diagram representing the forces acting on the combined human-exoskeleton system.
@@
We model the ankle exoskeleton as massless device, worn bilaterally, that only generates torque in the sagittal plane. We assume that the the human can resist the torque generated by the device if necessary. Thus the exoskeleton is represented by an ideal torque actuator at the ankle, and the net torque at the ankle is the sum of the human and the exoskeleton torques.
Rather than including the human torque $\tau_{\text{human}}$ as an ideal (instantaneous) torque actuator, we use a more realistic simplified actuation model where the direct human input to the system is the desired rate of torque development, $\dot{\tau}_{\text{human}}^{\text{des}}$. The stabilizable configurations are thus represented in joint angle-velocity-torque space, where the torque state represents the human torque only. Denoting the state space as $X\subset\mathbb{R}^{3}$, with $x_{1}=\theta,x_{2}=\dot{\theta}$ and $x_{3}=\tau_{\text{human}}$, and the input space as $U\subset\mathbb{R}$ with input $u=\dot{\tau}_{\text{human}}^{\text{des}}$, in state space form, our dynamics are
@@
The constants $g$, $m$, and $l$ represent the gravitational acceleration, mass of the body link, and leg length. The parameter $b$ represents minimal damping that is included make the dynamics less stiff.
We can thus limit the torque that the human can produce (MT) and the maximum rate of torque change (MRTD) of the model by constraining state $x_{3}$ and input $u$:
@@
Maximum plantar flexion torque $\tau_{\text{human}}^{\text{max pf}}$ is negative in our angle convention, while $\tau_{\text{human}}^{\text{max df}}$ denotes positive dorsiflexion torque. Lower magnitude of torque bounds correspond to weaker ankle strength, while lower bounds on $u$ correspond to reduced rate of torque development, as seen in older adults [8].
As, shown in Figure 2, the foot contact with the ground is determined by the normal force $F_{\text{normal}}$, the force due to friction $F_{\text{friction}}$ with coefficient of static friction $\mu$, and the location of the center of pressure (CoP). The CoP should remain within the base of support (BoS), which corresponds to the length of the foot. Full ground contact is enforced by three zero-moment point (ZMP) safety constraints which prevent the foot from tipping or slipping [28]:
@@
A human user may respond to the assistive torque of the exoskeleton in a range of ways. They may actively resist the torque generated by the device, or they may adapt their torque production to allow the exoskeleton to assist as much as possible. Studies with able-bodied participants, such as [19] have shown that young adults adapt quickly to an exoskeleton using a state-feedback controller, with minimal alteration to CoM kinematics. Thus we make the following assumptions related to the human-exoskeleton interaction:
@@
    The exoskeleton employs a state feedback control strategy and is unaware of the foot-ground interaction constraints- cannot reduce or increase torque production if a foot-ground interaction constraint is in danger of being violated
@@
    The human increases or decreases their torque production to account for the torque produced by the exoskeleton, in an effort to be optimally safe
@@
for $k\in\{1,2,3\}$. As we assume the exoskeleton torque is a function of state, these are actually state constraints on the human torque state $x_{3}$. Appendix 7 includes a full derivation of the equations of motion and foot-ground contact constraints.
@@ 4 Methods
We assess the effect of ankle exoskeletons on feasible stability by computing the BRS of the human-exo system for two exoskeleton controllers, over a range of sex- and age-adjusted model parameters.
@@ 4.1 HJB Reachability
The BRS can be computed using Hamilton Jacobi reachability tools. Hamilton Jacobi Bellman reachability uses HJB partial differential equations to compute implicit set representations of the backward reachable set [29]. We provide a brief overview of the method here, but we note that our analysis can be done with any nonlinear continuous time backward reachability tool. A more rigorous development of HJB and its use for solving reachability problems can be found in [30].
Given a dynamical system of the form (1) and a target set $S\subset X$, let $g(x)$ be a function whose zero sublevel set represents $S$, i.e., $S=\{x\mid g(x)\leq 0\}$. Typically $g(x)$ is a signed distance function. HJB reachability tools compute the backward reachable set by solving the PDE
@@
We assume that $u\in U\subset\mathbb{R}^{m}$ and that $U$ is a hyperrectangle such that $u$ is bounded elementwise, i.e. $u_{i}\in[\underline{u_{i}},\overline{u_{i}}]$. In this case, an analytical solution to Problem (7) is quite easy to find. As the optimization is over $u$, portions of the resulting expression that do not contain the control input can be ignored, and the problem reduces to
@@
The optimization problem is evaluated at a fixed time $t$ and state $x$, which means that each coefficient $\frac{\partial V(t,x)}{\partial x_{i}}\cdot f^{u}_{i}(x)$ is a constant. The solution can thus be found easily by letting $u_{i}=\underline{u_{i}}$ if the coefficient is positive and $u_{i}=\overline{u_{i}}$ otherwise. For system (3), equation (10) reduces to

### v3

@@ Abstract
Humans rely on ankle torque to maintain standing balance, particularly in the presence of small to moderate perturbations. Reductions in maximum torque (MT) production and maximum rate of torque development (MRTD) occur at the ankle with age, diminishing stability. Ankle exoskeletons are powered orthotic devices that may assist older adults by compensating for reduced torque and power production capabilities. They may also be able to assist with ankle strategies used for balance. However, the effect of such devices on standing balance in older adults is not well understood. Here, we model the effects ankle exoskeletons have on stability in physics-based models of healthy young and old adults, focusing on the potential to mitigate age-related deficits in MT and MRTD. Using backward reachability, a mathematical technique for analyzing the behavior of dynamical systems, we compute the set of stable center of mass positions and velocities for sex and age adjusted models of human standing balance with an ankle exoskeleton. We show that an ankle exoskeleton moderately reduces feasible stability boundaries in users who have full ankle strength. For individuals with age-related deficits, there is a trade-off. While exoskeletons augment stability at low center of mass velocities, they reduce stability in some high velocity conditions. Our results suggest that well-established control strategies must still be experimentally validated in older adults.
@@ 1 Introduction
Ankle exoskeletons are powered orthotic devices that can assist people with mobility impairments, such as older adults, by compensating for reduced muscle torque and power capabilities. Typically, they provide actuation about the ankle joint only in the sagittal plane, generating a push-off torque during dynamic activities such as walking and stair climbing. Exoskeletons have been shown to reduce metabolic cost of transport during walking in both younger [1] and older adults [2]. The main appeal of lower-limb exoskeletons to older adults, however, is their perceived potential to reduce fall risk [3]. It is therefore crucial to develop our understanding of the effects of ankle exoskeleton assistance on stability, which is currently incomplete.
@@
Torque production at the ankle is an important contributor to balance, which suggests that ankle exoskeletons have the potential to improve stability during standing. Unimpaired adults tend to rely on modulating ankle torques to maintain a stable standing position, only switching to a hip strategy in response to large perturbations [4]. In older adults, maximum available plantar flexion and dorsiflexion torques are lower than in younger adults, reducing the range of perturbations that can be accommodated with an ankle strategy [5]. This reduction in maximum ankle torque (MT) has been shown in multiple studies to be correlated with decreased performance in balancing tasks [5] and a higher risk of falls [6, 7]. The rate at which this torque can be produced also declines with age [8], although the effect of a lower maximum rate of torque development (MRTD) on standing balance is inconclusive. Reduced MRTD at the knee is associated with a history of falls [9, 10], while plantar flexor MRTD is significantly correlated with lower performance on a single standing leg balance test for older men, but not for older women [11]. Other studies have found that lower ankle MRTD is not a strong predictor of fall likelihood [12, 13]. Still, lower MRTD likely contributes, independently of lower MT, to increased fall risk in older adults [13].
@@
Although ankle exoskeletons can, in principle, mitigate both MT and MRTD deficits by producing large amounts of torque quickly, the effect they have on stability is unclear. Depending on the population, device, controller, and perturbation paradigm, ankle exoskeletons have been reported to enhance stability [14, 15, 16], slightly reduce stability [17, 15, 18], or have no effects on stability [19, 20]. Notably, the majority of balance studies for powered exoskeletons have focused on young and able-bodied users, who do not have impaired stability. As these devices are likely to be marketed to enhance mobility in older populations [3], it is imperative to understand how they may affect the stability of users with age-related reduction in MT and MRTD. We currently do not know how these age-related changes may interact with additional exoskeleton assistance at the ankle. Because perturbation experiments on older adults are expensive, time-consuming, and risky, developing a generalizable model-based characterization of standing balance with ankle exoskeletons is a valuable, complementary approach. Modeling can provide insights into the mechanisms underlying exoskeleton-assisted standing balance, including how constraints on the human-exoskeleton system are affected both by the device and by age-related joint-level torque deficits.
@@
Bounds on human stability for feet-in-place activities such as standing may be characterized by computing the set of all initial body center of mass (CoM) positions and velocities from which it is possible to stabilize to quiet standing, meaning that the final CoM position is above the foot and the CoM velocity is close to zero [21]. This set is sometimes called the feasible stability region [22, 23].For a simplified linear inverted pendulum these are the well-known extrapolated center of mass boundaries (XCoM) [24]. Herein we will use the term stabilizable region, abbreviated ‘SR’ [25].
@@
Stabilizable regions have been computed for various legged dynamical systems [26, 27], including one example where the boundary was partially computed for a model of a human wearing an optimally controlled hip-knee-ankle exoskeleton [28]. The full stabilizable region under aging-related biological constraints and different ankle exoskeleton controllers has not been explored. Furthermore, formal stability properties of this region have not been defined. Invariance is of particular interest. If the SR is controlled invariant, this means that there always exists a controller that allows the human to stay within the region. If it is not, then there may be states within this region from which the human can pass through the SR, but not stabilize. Invariance guarantees that the SR does not include states which merely pass through the set. This property is especially important to verify when there are constraints on the control input. Such constraints may correspond to biologically meaningful quantities such as maximum joint torque, or to environmental constraints such as the frictional characteristics of the support surface.
@@
We previously introduced a framework for determining the complete SR boundary for human dynamical systems [25]. Our approach formulates the SR boundary as the solution of a single partial differential equation rather than a brute-force algorithmic search as in prior work, e.g. [26, 28]. Moreover we provide strong formal invariance and stability guarantees over the entire SR in the presence of both biological and environmental constraints.
@@
Here we present a generalizable, model-based understanding of how an ankle exoskeleton may alter the domain of invariant stable motions available to a user in the presence of joint level functional changes associated with aging, including reduced MT production and MRTD. To our knowledge, this is the first study that attempts to systematically understand the combined effect of ankle exoskeleton assistance and age-related ankle torque deficits on feasible stability. Using common exoskeleton control strategies, we analyze the effect that exoskeletons have on the invariant stabilizable region for models of young and old adults. We show that while exoskeletons can indeed enhance ankle torque production and improve stability under certain conditions, they can also act as a disturbance, hindering stability in safety-critical regions of the state space.
@@ 2 Preliminaries
Our methods for computing stabilizable regions (SRs) are based on two notions, controlled invariance and backward reachability, which we formally define in this section.
@@ 2.1 Controlled invariance and backward reachability
Figure 1: Illustration of a target set (green) and the set of all states from which it can be reached. This set is called the backward reachable set (BRS, in purple). The state $\hat{x}_{0}$ is in the BRS, because there exists at least one controller, $u_{1}$, that drives the system to the target set. No such controller exists for $\tilde{x}_{0}$ so it is not in the BRS. If the target set is controlled invariant then the union of the target set and the BRS is also invariant. Here, the BRS was computed with model parameters corresponding to a young male, using our method described in Section 4. The gray region is computed using the extrapolated center of mass (XCoM) method from [24], which we have converted here to angular coordinates.
@@
A set $\Omega\subset X$ is controlled invariant if, for any initial condition in the set, there exists a controller that maintains the system state within $\Omega$:
@@
The BRS contains states from which the system can reach the target set at precisely time $T$. The set of states from which the target set can be reached at $t\leq T$ is called the backward reachable tube, and is a union of backward reachable sets:
@@ 2.2 Relating invariance and reachability to human stability
The relationship between human stability and our mathematical definitions is illustrated in Figure 1. Assume that the dynamics (1) represent a parameterized model of a standing human. In the figure, the green ellipse represents a controlled invariant target set of states corresponding to quiet standing. For the purposes of the illustration the set is depicted to be larger than it is. The purple region is the true-to-size backward reachable set that we compute using a constrained dynamical model with parameters corresponding to a young male, as described later in Section 3. The boundaries of this set correspond to forward and backward fall boundaries, and to constraint failures. Outside of these boundaries the ankle strategy fails and the model must take some action to avoid a fall, such as taking a step. For visual clarity we only depict the portion of the state space corresponding to forward angular velocities. Because the target set is controlled invariant, the resulting purple safe region is as well. For comparison, we show the same bounds computed using the XCoM from [24], in light gray. XCoM is often used by biomechanists to estimate dynamic stability boundaries for humans as they stand and walk [29].
To determine the stability bounds of the human-exoskeleton system, we compute reachable sets with the dynamics of an ankle exoskeleton controller added to (1), the details of which are presented in Section 3 and 4. We construct invariant target sets corresponding to quiet standing with exoskeleton assistance, and then compute the backward reachable sets. The resulting sets then delineate the feasible stabilizable boundaries with ankle exoskeleton assistance.
@@ 3 Human-Exoskeleton Model
Figure 2: Free body diagram of the combined human-exoskeleton system.
@@
We model the ankle exoskeleton as a massless device, worn bilaterally, that only generates torque in the sagittal plane. We assume that the human can resist the torque generated by the device if necessary. Thus the exoskeleton is represented by an ideal torque actuator at the ankle, and the net torque at the ankle is the sum of the human and the exoskeleton torques.
Rather than including the human torque $\tau_{\text{human}}$ as an ideal (instantaneous) torque actuator, we use an actuation model where the direct human input to the system is the desired rate of torque development, $\dot{\tau}_{\text{human}}^{\text{des}}$. The stabilizable configurations are thus represented in joint angle-velocity-torque space, where the torque state represents the human torque only. Denoting the state space as $X\subset\mathbb{R}^{3}$, with $x_{1}=\theta,x_{2}=\dot{\theta}$ and $x_{3}=\tau_{\text{human}}$, and the input space as $U\subset\mathbb{R}$ with input $u=\dot{\tau}_{\text{human}}^{\text{des}}$, in state space form, our dynamics are
@@
The constants $g$, $m$, and $l$ represent the gravitational acceleration, mass of the body link, and leg length. The parameter $b$ represents minimal damping that is included to make the dynamics less stiff.
We can thus limit the human MT and MRTD of the model by constraining state $x_{3}$ and input $u$:
@@
Maximum plantar flexion torque $\tau_{\text{human}}^{\text{max pf}}$ is negative in our angle convention, while $\tau_{\text{human}}^{\text{max df}}$ denotes positive dorsiflexion torque. Lower magnitudes of torque bounds correspond to weaker ankle strength, while lower bounds on $u$ correspond to reduced rate of torque development, as observed in older adults [8].
Foot contact with the ground is determined by the normal force $F_{\text{normal}}$, frictional force $F_{\text{friction}}$ with coefficient of static friction $\mu$, and the location of the center of pressure (CoP). The CoP should remain within the base of support (BoS), which corresponds to the length of the foot. Full ground contact is enforced by three zero-moment point (ZMP) safety constraints which prevent the foot from tipping or slipping [30]:
@@
A human user could respond to the assistive torque of the exoskeleton in a range of ways. They may actively resist the torque generated by the device, or they may adapt their torque production to allow the exoskeleton to assist as much as possible. Studies with able-bodied participants, such as [20] have shown that young adults adapt quickly to an exoskeleton using a state-feedback controller, with minimal alteration to CoM kinematics. Thus we make the following assumptions related to the human-exoskeleton interaction:
@@
    The exoskeleton employs a state feedback control strategy and is unaware of the foot-ground interaction constraints. It cannot reduce or increase torque production if a constraint is in danger of being violated
@@
    The human adapts their torque production to account for the torque produced by the exoskeleton
@@
for $k\in\{1,2,3\}$. As we assume the exoskeleton torque is a function of state, these are actually state constraints on the human torque state $x_{3}$. Appendix I of the supplementary material includes a full derivation of the equations of motion and foot-ground contact constraints.
@@ 4 Methods
We assess the effect of ankle exoskeletons on feasible stability by computing the BRS of the human-exoskeleton system for two exoskeleton controllers, over a range of sex- and age-adjusted model parameters.
@@ 4.1 HJB Reachability
The BRS can be computed using Hamilton Jacobi reachability tools. Hamilton Jacobi Bellman reachability uses HJB partial differential equations to compute implicit set representations of the backward reachable set [31]. We provide a brief overview of the method here, but we note that our analysis can be done with any nonlinear continuous time backward reachability tool. A more rigorous development of HJB and its use for solving reachability problems can be found in [32]. Given a dynamical system of the form (1) and a target set $S\subset X$, let $g(x)$ be a function whose zero sublevel set represents $S$, i.e., $S=\{x\mid g(x)\leq 0\}$. Typically $g(x)$ is a signed distance function. HJB reachability tools compute the backward reachable set by solving the PDE
@@
We assume that $u\in U\subset\mathbb{R}^{m}$ and that $U$ is a hyperrectangle such that $u$ is bounded elementwise, i.e. $u_{i}\in[\underline{u_{i}},\overline{u_{i}}]$. In this case, a simple analytical solution to Problem (7) exists. As the optimization is over $u$, portions of the resulting expression that do not contain the control input can be ignored, and the problem reduces to
@@
The optimization problem is evaluated at a fixed time $t$ and state $x$, which means that each coefficient $\frac{\partial V(t,x)}{\partial x_{i}}\cdot f^{u}_{i}(x)$ is a constant. The solution can thus be found by letting $u_{i}=\underline{u_{i}}$ if the coefficient is positive and $u_{i}=\overline{u_{i}}$ otherwise. For system (3), equation (10) reduces to